OpenStudy (anonymous):

-a2 - 3b3 + c2 + 2b3 - c2 if a = 3, b = 2, and c = -3.

OpenStudy (anonymous):

substitute a=3, b=2 and c=-3 into the equation

OpenStudy (anonymous):

just replacing your a value with 3, b with 2 and c with -3 for example y = 3x^2 if x =2 y = 3(2)^2 y = 3 x 4 = 12

OpenStudy (anonymous):

i keep getting the wrong answer though